Manager of a monopoly inverse demand curve of P = 270 - 20Q and C= 5 + 30Q Max profits are :

715

755

751

725
Answer #1

Answer

The correct answer is (a) 715

Profit(Pr) = TR - C

where TR = Total Revenue= PQ = (270 - 20Q)*Q and C = Cost = 5 + 30Q

=> Pr = (270 - 20Q)*Q - (5 + 30Q) -------------------Profit Function

Maximize : Pr = (270 - 20Q)*Q - (5 + 30Q)

First Order condition:

d(Pr)/dQ = 0 => 270 - 40Q - 30 = 0

=> Q = 6

Putting this in Profit function we get:

Profit(Pr) = (270 - 20*6)*6 - (5 + 30*6) = 715

Hence, Max Profits are 715.

Hence, the correct answer is (a) 715